The temperature of an object is given by T ( t ) = 280 + 1.5 t 2 e − 0.12 t where t is measured in minutes and T is measured in Kelvins (abbreviated K). (a) Find the rate of change of the rate of change of temperature at the instant t = 20 minutes.
1
Expert's answer
2020-07-27T19:07:25-0400

Given


"T(t)=280+1.5t^2e^{-0.12t}"

"T'(t)=(280+1.5t^2e^{-0.12t})'="

"=1.5(2t)e^{-0.12t}+1.5t^2(-0.12e^{-0.12t})="

"=3te^{-0.12t}-0.18t^2e^{-0.12t}"

"T''(t)=(3te^{-0.12t}-0.18t^2e^{-0.12t})'="

"=3e^{-0.12t}+3t(-0.12e^{-0.12t})-"

"-0.18(2t)e^{-0.12t}-0.18t^2(-0.12e^{-0.12t})="

"=3e^{-0.12t}-0.72te^{-0.12t}+0.0216t^2e^{-0.12t}"


"T''(20)=3e^{-0.12(20)}-0.72(20)e^{-0.12(20)}+"

"+0.0216(20)^2e^{-0.12(20)}\approx-0.25K/min^2"

The rate of change of the rate of change of temperature at the instant t = 20 minutes is "-0.25\ K/min^2."